Consider a monopolistic competitor who has the following demand function p=140-4Q.In additionsuppose that his total cost function is given Tc=10+5Q2 . Find his equilibrium price and quantity


1
Expert's answer
2021-11-14T17:00:44-0500

demand function p=140-4Q

total cost function is given Tc=10+5Q


MR"=" MC;

MC"=10Q;"

"MR=((140-4Q)\u00d7Q )'=(140Q-4Q^2)'=140-8Q;"

"10Q=140-8Q; \n18Q=140;"

"Q=" "7.8;"

"P=140-8\u00d77.8; P=140-31.2;"

"P=" $"108.8;"